OpenStudy (anonymous):

Factor 16x^2 + 49. Check your work.

OpenStudy (anonymous):

This is what i have so far:

OpenStudy (anonymous):

16x2 + 49 = 16x2 – (-49) = (4x+7i)(4x-7i)

OpenStudy (anonymous):

but when I checked my work, the result was 16x^2 - 49... so is that still right?

OpenStudy (anonymous):

youre doing it right, your checkin is wrong. i x i = -1

OpenStudy (anonymous):

so I just substitute all the i's with -1's and then multiply?

OpenStudy (anonymous):

no i into i is -1

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

what do u mean "into"?

OpenStudy (anonymous):

i multiplied by i

OpenStudy (anonymous):

so i fix the last part right? -49i^2

OpenStudy (anonymous):

yes - 49i^2 = -49 x -1 =49

OpenStudy (anonymous):

ok so,

OpenStudy (anonymous):

16x^2 +28x -28x +49i^2 16x^2 + 49i^2

OpenStudy (anonymous):

bingo!

OpenStudy (anonymous):

oh and i have one more

OpenStudy (anonymous):

x^2 + 64

OpenStudy (anonymous):

this is what i got:

OpenStudy (anonymous):

x^2 – (-64) (x+ 8i)(x- 8i)

OpenStudy (anonymous):

so now i have to check it

OpenStudy (anonymous):

so it's exactly the same right?

OpenStudy (anonymous):

x^2 - 64i^2

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

wait so that's what u get when u check it?

OpenStudy (anonymous):

so its: x^2 – 64i^2 x^2 – 64(-1)^2 x^2 +64

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

no prob